# Information Sifting Funnel: Privacy-preserving Collaborative Inference Against Model Inversion Attacks

### v1

@@ 2.3. Related defense work
To protect the input data collected by edge devices in collaborative inference from potential leakage, He et al. (He et al., 2019, 2020) highlighted the necessity of developing defense strategies while identifying these threats. Building on a detailed analysis of attack methods, subsequent defense techniques have gradually evolved in recent years. These defenses can be broadly categorized into three types: perturbation-based defenses, information bottleneck-based defenses, and defenses combining neural network depth with information bottleneck principles. The following sections provide a detailed overview of each of these defense strategies.
Perturbation-based defense. This approach introduces carefully designed noise perturbations into the gradients and output features of the edge model during training to reduce the likelihood of adversaries extracting meaningful information and defend against MIA. He et al. (He et al., 2020) initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al. (Wang et al., 2022) extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features. Although these methods enhance defenses, they significantly impact model usability, often reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA. This is because sufficiently expressive decoders in attack models can adapt to small noise or feature dropout, leveraging reconstruction capabilities similar to denoising auto-encoders (DAE) (Gondara, 2016), potentially increasing attack robustness. In centralized MIA defenses, Struppek et al. (Struppek et al., 2023) proposed perturbing training labels by setting the label smoothing factor to a negative value, which significantly reduced the effectiveness of MLE-based MIA and offers an alternative perspective for designing defense strategies.
IB-based defense. This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between features and outputs $y$, thereby maintaining predictive performance (Wang et al., 2021). The objective of this defense can be formally expressed as follows:
@@
Initially, Wang et al. (Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for collaborative inference and required substantial computational resources when applied to intermediate features. To address this, Peng et al. (Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which optimizes the correlation between inputs, features, and outputs using dependency algorithms. BiDO reduces the correlation between inputs and features while strengthening the correlation between features and outputs, balancing defense and performance. Additionally, Wang et al. (Wang et al., 2024) and Duan et al. (Duan et al., 2023) explored direct mutual information estimation between inputs and features using methods based on mutual information definitions and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the optimization objective. IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research. However, these methods face limitations: BiDO relies on kernel functions to constrain nonlinear correlations, and its performance is highly sensitive to hyper-parameter choices, such as kernel bandwidth. Poor parameter selection can result in suboptimal optimization. Methods based on mutual information definitions are constrained by sample scale to estimate $I(x,z)$.
@@
NND & IB-based defenses. Building on the natural advantages of deep neural networks in resisting inference attacks and incorporating information bottleneck theory, this approach designs targeted loss functions to further enhance defense effectiveness. Building on this, Ding et al. (Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict the transmission of redundant information. Subsequently, Azizian et al. (Azizian and Bajić, 2024) proposed an improved defense method by introducing encoders at the edge and decoders at the cloud, forming an auto-encoder (AE) structure. The encoder at the edge increases network depth and reduces feature dimensionality, achieving the information bottleneck effect, while the decoder at the cloud restores information to maintain predictive performance. Additionally, an $l_{1}$-norm constraint constraint is applied during feature transmission to reduce attack risks, and adversarial training is used to enhance defense resilience against inference attacks.
@@
However, this class of methods has certain limitations. First, increasing the depth of neural networks and adding additional modules inevitably raises computational overhead and hardware requirements on edge devices, limiting the method’s applicability. Second, the introduction of adversarial training not only obscures certain aspects of the original defense strategy’s effectiveness but also introduces additional challenges. It is less effective against future, more advanced attacks, increases the difficulty of fine-tuning edge models, and may degrade inference performance (Zhang et al., 2019). Additionally, adversaries can adapt their attack models using similar adversarial strategies, reducing the overall effectiveness of the defense (Jin et al., 2024). Thus, while these methods enhance defense performance, their implementation requires careful consideration of computational resources, model adaptability, and suitability for specific application scenarios.
@@ 3.1. Knowledge background of adversary
    Black-box Scenario: The adversary has only black-box access to the model, restricting their capability to observing the intermediate features. Additionally, their understanding of the edge model’s architecture is restricted to prior knowledge or incomplete assumptions.
@@
    Gray-box Scenario: The adversary has partial knowledge of the edge model’s architecture but remains unaware of defensive modifications, such as increased depth or additional hidden layers.
The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their inversion model. For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, the adversary can only construct an approximate attack model based on partial knowledge of the initial layers. However, in the gray-box scenario, the adversary can address this limitation by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their inversion model.
Knowledge shared by adversaries in different scenarios. First, in attack scenarios, the adversary is assumed to understand the edge model’s core task. To enhance their capability, the attack model in Gen-based MIA is trained on data distributions closely matching the target distribution, as detailed in Section 6. Second, the adversary has access to lossless transmission features and their dimensional distribution, which are essential for attacks, as the features represent high-dimensional input data and the distribution informs the design of the attack model.
This study focuses on image classification tasks, though the proposed defense methods extend to other tasks where the edge model primarily performs feature extraction. Tasks like segmentation, detection, and classification rely on convolutional layers for shallow feature extraction, with task-specific variations having minimal impact on reconstruction objectives. Prior work (Azizian and Bajić, 2024) has also demonstrated input reconstruction effectiveness in object detection using Gen-based MIA.
Moreover, image data, due to its high dimensionality, is especially vulnerable to MIA attacks, as shallow convolutional layers retain redundant information (Zhang et al., 2020; He et al., 2016). This makes defenses in image-based tasks more demonstrable. In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in existing researches for achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022). This work focuses on image data to explore privacy-preserving techniques, leveraging its interpretability and the intuitive nature of defense effectiveness to validate the proposed methods.
@@ 3.2. Attack strategy of the adversary
In the white-box attack scenario, the adversary can perform attacks using the rMLE-based MIA method proposed in (He et al., 2019, 2020). To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
@@ 4. Theoretical Analysis of MIA in CI
This section analyzes key factors affecting MIA effectiveness in collaborative inference, introduces a formulaic representation of its principles, and establishes a difficulty criterion from an information-theoretic perspective. We validate this criterion through experiments and use it to guide the design of SiftFunnel.
@@ 4.1. Principles analysis of MIA
MIAs are categorized into MLE-based and Gen-based methods, both aiming to reconstruct input data using the target model’s output. MLE-based MIA calculates the loss between transmitted features and the edge model’s output, then iteratively applies gradient descent to optimize the input, achieving reconstruction of the original data. The attack principle is formulated as follows:
@@
According to (He et al., 2019), the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA. The method relies on gradient descent, where the loss values depend on output features and gradients are computed using the target model’s parameters.
@@
[…]
@@
Where $x^{\prime}$ represents samples from an auxiliary data distribution $p\left(x^{\prime}\right)$. The loss function, typically MSE, measures the distance between $x^{\prime}$ and the reconstructed data $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ or $z=f_{\text{edge }}\left(x^{\prime}\right)$. The trained generator $G$ must achieve the desired mapping $G\approx{f_{\text{edge}}}^{-1}$ to reconstruct $x^{*}$.
@@
The rationale is that if the information in $f_{\text{edge }}\left(x^{\prime}\right)$ is insufficient or cannot independently represent $x^{\prime}$’s features, the loss in (4) will struggle to converge as $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ remains ambiguous, limiting its training. These conditions must be met simultaneously. For instance, even with random feature drops, if the remaining features effectively distinguish inputs, the attack can still achieve satisfactory reconstruction (He et al., 2020). Such frameworks resemble DAE (Gondara, 2016), where the decoder can adapt to noise or feature drops.
Gen-based MIA demonstrates superior reconstruction performance compared to MLE-based MIA in defense scenarios. First, it avoids relying on model parameters for gradient computation, eliminating challenges from non-convex optimization. Second, it depends only on edge model output, offering greater flexibility. Furthermore, existing defenses struggle to effectively remove redundant feature information, allowing Gen-based MIA to overcome defenses if the generator is capable of learning the edge model’s reverse mapping.
@@ 4.2. Difficulty criterion for MIA in CI
Above, we summarized the factors influencing MIA effectiveness through a formulaic representation of its principles. However, this alone does not provide a concrete basis for evaluating the difficulty of implementing MIA. To address this, we will formalize and analyze these factors from an information-theoretic perspective to establish a clear criterion.
@@
From Equation (5), we observe that complex mappings and information loss during neural network transmission reduce $I(x;z)$ and increase $H(x|z)$, making non-convex optimization more difficult. Consequently, this increases the challenge for MLE-based MIA to reconstruct $x$ from $z$.
An inappropriate loss function or insufficient output information can further increase $H(x|z)$ and decrease $H(z)$. It is important to note that a low $\delta(z)$ does not necessarily imply a low $H(z)$. For instance, even if 80% of $z$’s elements are randomly dropped, $H(z)$ could remain high, as small amounts of information can still exhibit high entropy (He et al., 2020). Therefore, the performance of MLE-based MIA is directly influenced by $H(x|z)$, $H(z)$, and $\delta(z)$.
To further illustrate the influence of mutual information and entropy on the difficulty of MIA, we can derive the following relationship for any $x\to z\to x^{*}$ using Fano’s inequality (Arrow, 1969; Xu et al., 2024):
@@
Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore inequality (6) can be written as
@@
Since $H(x)$ and $|\mathbb{X}|$ are influenced by training data, they are typically assumed to be objective and invariant. Therefore, through inequality (7), we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x|z)$, $H(z)$, and $I(x,z)$.
From Equation (4), Gen-based MIA requires sufficient input information $\delta(z)$ and high $H(z)$ to ensure maximum separability and enable loss convergence, allowing the generator to simulate reverse mapping effectively. Notably, $\delta(z)$ and $H(z)$ are not directly proportional, as high average information does not always guarantee separability. A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$. The relationship in Inequality (7) is also applicable here.
The validity of related research can also be explained by both. For example, (Yang et al., 2019) reduces $\delta(z)$ by pruning output information, disrupting training and reconstruction, while (Zhang et al., 2023) increases entropy via nonlinear functions to enhance generator performance. Thus, $H(z)$ and $\delta(z)$ are critical factors influencing Gen-based MIA.
The analysis indicates that the difficulty of implementing both types of attacks is inversely proportional to mutual information, entropy, and the amount of information. The difficulty of MIA can be formulated as follows, where $k_{1}$ and $k_{2}$ are proportionality constants.
@@
Based on this criterion, existing defenses (Wang et al., 2021; Peng et al., 2022; Duan et al., 2023) using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA. Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x|z)$. The defensive effect of deeper neural networks (Ding et al., 2024; Azizian and Bajić, 2024) can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$.
This is primarily because, without skip connections (Hao et al., 2025), mutual information decreases progressively during the forward propagation in neural networks. In contrast, residual networks with skip connections add the output of one layer to its input, which may prevent $\delta(z)$ from decreasing and, in some cases, even cause it to increase.
4.3. Experimental Demonstration of the Criterion and Guidance for Defense Design
To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces spatial size while increasing channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block. We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of mutual information. The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan representation (DV bound) (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.

### v4

@@ 2.3. Related defense work
To safeguard input data collected by edge devices in CI from potential leakage, He et al. (He et al., 2019, 2020) emphasized the importance of developing effective defense strategies while identifying these vulnerabilities. Existing defenses can be classified into three main categories: perturbation-based, IB-based, and neural network depth (NND) & IB-based approaches. The following sections outline these strategies.
Perturbation-based defense. This approach introduces carefully designed noise perturbations into the gradients for training or trained edge model output features to reduce the likelihood of adversaries extracting meaningful information and defending against MIA. He et al. (He et al., 2020) initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al. (Wang et al., 2022) extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features. Although these methods enhance defenses, they significantly impact model usability, typically reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA. The root cause is that the decoder in Gen-based MIAs is sufficiently expressive to adapt to noise or feature dropout, similar to the training of a denoising auto-encoder (DAE) (Gondara, 2016), potentially increasing attack robustness. In centralized MIA defenses, Struppek et al. (Struppek et al., 2023) proposed perturbing training labels with a negative label smoothing factor, significantly reducing MIA effectiveness and offering a novel perspective for designing defense strategies.
IB-based defense. This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between $z$ and outputs $y$, thus maintaining inference performance (Wang et al., 2021). The objective of this defense can be formally expressed as follows:
@@
Initially, Wang et al. (Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for CI and required substantial computational resources for intermediate feature processing. To address this, Peng et al. (Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which uses dependency algorithms to optimize correlations among inputs, features, and outputs. BiDO reduces input-feature correlation while enhancing feature-output correlation, achieving a balance between defense and performance. Additionally, Wang et al. (Wang et al., 2024) and Duan et al. (Duan et al., 2023) explored mutual information estimation between $x$ and $z$ using methods based on mutual information definition and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the training loss function. IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research. However, these methods face limitations: BiDO relies on kernel functions for non-linear correlation constraints, with performance highly sensitive to hyper-parameters like kernel bandwidth; poor parameter selection can lead to suboptimal results. Methods based on mutual information estimation are constrained by sample scale to compute.
@@
NND & IB-based defense. Leveraging deeper networks and IB theory, this approach enhances defense effectiveness with targeted loss functions. Ding et al. (Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict redundant information transmission. Subsequently, Azizian et al. (Azizian and Bajić, 2024) further improved defenses by introducing an auto-encoder (AE) structure, with encoders at the edge and decoders in the cloud. The edge encoder increases network depth and reduces feature dimensionality to achieve the IB effect, while the cloud decoder restores information to maintain predictive performance while applying an $l_{1}$-norm constraint to reduce attack risks. Adversarial training (Ding et al., 2024) is also incorporated to strengthen robustness against MIAs.
@@
However, this method has certain limitations. Increasing network depth and adding modules raise computational and hardware demands on edge devices, limiting applicability. Adversarial training, while strengthening defenses, introduces additional challenges—it is less effective against advanced attacks, complicates fine-tuning, and degrades inference performance (Zhang et al., 2019). Additionally, adversaries can adapt their attack models using similar adversarial strategies, weakening defense effectiveness (Jin et al., 2024). While these methods improve defenses, their implementation requires careful consideration of computational resources, model adaptability, and application suitability. SiftFunnel enhances computational efficiency by deepening edge model structures and adopting a lightweight design, exhibiting IB-like characteristics after training. Thus, our method falls under the NND & IB-based defense strategy.
@@ 3.1. Knowledge background of adversary
    Black-box Scenario: The adversary has only black-box access to the model, restricting their capability to observe the intermediate features. Their understanding of the edge model’s architecture is also restricted to prior knowledge or incomplete assumptions.
@@
    Gray-box Scenario: The adversary has partial knowledge of the edge model’s architecture but remains unaware of defensive modifications, such as increased hidden layers.
The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their attack model. For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, they can only construct an approximate attack model based on partial knowledge of the initial three layers. However, adversaries can address this limitation in this scenario by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their attack model.
Knowledge shared by adversaries in different scenarios. The adversary is assumed to understand the edge model’s task and enhances their attack by training Gen-based MIA on a closely matching data distribution (Section 6.1). Additionally, the adversary has access to lossless transmission features and their dimensional distribution, which facilitate optimizing the MIA loss function and training the attack model.
This study focuses on image classification but emphasizes that the defense methods apply to other tasks where the edge model primarily handles feature extraction. Since segmentation, detection, and classification also rely on edge-processed features, MIA and defenses are task-agnostic. Prior work (Azizian and Bajić, 2024) has demonstrated input reconstruction and defense effectiveness in object detection using Gen-based MIA.
Moreover, image data is particularly vulnerable to MIAs due to its high dimensionality, as shallow convolutional layers retain more redundant information (Zhang et al., 2020; He et al., 2016). This increases the challenge of protecting image data, making it an ideal test case for privacy defenses. In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022). This study focuses on image data due to its interpretability, making it suitable for evaluating privacy defenses.
@@ 3.2. Attack strategy of the adversary
In the white-box attack scenario, the adversary can perform attacks using the MLE-based MIA method proposed in (He et al., 2019, 2020). To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
@@ 4.1. Principles analysis of MIA
MIAs are categorized into MLE-based and Gen-based methods, both aiming to reconstruct input data using the output of the edge model. MLE-based MIA calculates the loss between transmitted features and the output, then iteratively applies gradient descent to optimize the input, achieving reconstruction of the original data. The attack principle is formulated as follows:
@@
According to (He et al., 2019), the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA. The method relies on gradient descent, where the loss values depend on output features, and gradients are computed using the target model’s parameters.
@@
When dealing with complex edge model mappings or models protected by defense techniques, directly computing gradients in MLE-based MIA can significantly hinder optimization performance. To address this, He et al. (He et al., 2019, 2020), building on the method from (Yang et al., 2019), developed a reverse mapping network for the edge model, which is categorized as Gen-based MIA. The key idea is to use a generator to approximate the reverse mapping of the edge model. Different methods adopt various strategies to train the generator. For instance, Yang et al. (Yang et al., 2019) trained a decoder using auxiliary task-related datasets, while Yin et al. (Yin et al., 2023) utilized intercepted feature data for training. The principle of this attack is summarized as follows.
@@
Where $x^{\prime}$ represents samples from an auxiliary data distribution $p\left(x^{\prime}\right)$. The loss function, typically mean square error, measures the distance between $x^{\prime}$ and the reconstructed data $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ or $z=f_{\text{edge }}\left(x^{\prime}\right)$. The trained generator $G$ must achieve the desired mapping $G\approx{f_{\text{edge}}}^{-1}$ to reconstruct $x^{*}$.
@@
The rationale is that if the information in $f_{\text{edge }}\left(x^{\prime}\right)$ is insufficient or cannot independently represent $x^{\prime}$’s features, the loss in (4) will struggle to converge as $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ remains ambiguous, limiting its training. Among them, the maximum separability is more decisive. For instance, even with random feature drops, if the remaining features distinguish inputs well, the attack can still reconstruct effectively (He et al., 2020). Such frameworks resemble DAE (Gondara, 2016), where the decoder can adapt to noise or feature drops.
Gen-based MIA demonstrates superior reconstruction performance compared to MLE-based MIA in defense scenarios. First, it avoids relying on model parameters for gradient computation, eliminating challenges from non-convex optimization. Second, it depends only on the output of the edge model, offering greater flexibility. Furthermore, existing defenses struggle to effectively remove redundant feature information, allowing Gen-based MIA to overcome defenses if the generator is capable of learning the edge model’s reverse mapping.
@@ 4.2. Difficulty criterion for MIA in CI
We summarized the factors influencing MIA effectiveness using a formulaic representation. To further evaluate the difficulty of MIA implementation, we will formalize and analyze these factors from an information-theoretic perspective to establish a clear criterion.
@@
Factors influencing MLE-based MIA. From Equation (5), we observe that complex mappings and information loss during neural network forward propagation reduce $I(x;z)$ and increase $H(x|z)$, making non-convex optimization with backpropagation more difficult. Consequently, reconstructing $x$ from $z$ becomes more difficult for MLE-based MIA.
These methods primarily compare $z$ with generated features, but a poorly trained or protected model can produce $z$ with high $\delta(z)$ yet low $H(z)$, indicating suboptimal separability. This weakens attack effectiveness despite the apparent information sufficiency. Thus, MLE-based MIA difficulty is more influenced by $I(x;z)$, $H(x|z)$, and feature entropy $H(z)$.
We can demonstrate the influence of mutual information and entropy on MLE-based MIA using the following relationship for any $x\to z\to x^{*}$, derived from Fano’s inequality (Arrow, 1969; Xu et al., 2025):
@@
Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore, inequality (6) can be written as
@@
Since $H(x)$ and $|\mathbb{X}|$ are influenced by training data, they are typically assumed to be objective and invariant. Therefore, we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x\mid z)$, $H(z)$, and $I(x,z)$.
Factors influencing Gen-based MIA. From Equation (4), Gen-based MIA requires sufficient input information $\delta(z)$ and high $H(z)$ to ensure maximum separability and enable loss convergence, allowing the generator to simulate reverse mapping effectively. A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$. At the same time, the relationship in Inequality (7) also holds here (Zhang et al., 2023). Therefore, in addition to the factors influencing MLE-based MIA, Gen-based MIA is further effectively affected by $\delta(z)$.
Criterion formula. The analysis indicates that the difficulty of implementing both types of MIA is inversely proportional to the mutual information, entropy, and output information. Therefore, the difficulty criterion for MIA can be formulated as follows, where $k_{1}$ and $k_{2}$ are proportionality constants:
@@
Based on this criterion, existing defenses (Wang et al., 2021; Peng et al., 2022; Duan et al., 2023) using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA. Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x\mid z)$. The defensive effect of deeper neural networks (Ding et al., 2024; Azizian and Bajić, 2024) can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$. This is mainly because, without skip connections (Hao et al., 2025), mutual information progressively decreases during forward propagation in neural networks. In contrast, residual networks with skip connections preserve or even enhance $\delta(z)$ by adding layer outputs to inputs, but the decline in mutual information still poses challenges for MIA (see section 4.3 for experimental cases).
4.3. Demonstration and Guidance for Defense
To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces the spatial size while increasing the channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block. We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of $I(x,z)$. The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan (DV) representation (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.